# Geometric quantum thermodynamics: A fibre bundle approach

### v2

@@ Abstract
Classical thermodynamics is a theory based on coarse-graining, meaning that the thermodynamic variables arise from discarding information related to the microscopic features of the system at hand. In quantum mechanics, however, where one has a high degree of control over microscopic systems, information theory plays an important role in describing the thermal properties of quantum systems. Recently, a new approach has been proposed in the form of a quantum thermodynamic gauge theory, where the notion of redundant information arises from a group of physically motivated gauge transformations called the thermodynamic group. In this work, we explore the geometrical structure of quantum thermodynamics. Particularly, we do so by explicitly constructing the relevant principal fibre bundle. We then show that there are two distinct (albeit related) geometric structures associated with the gauge theory of quantum thermodynamics. In this way, we express thermodynamics in the same mathematical (geometric) language as the fundamental theories of physics. Finally, we discuss how the geometric and topological properties of these structures may help explain fundamental properties of thermodynamics.
@@ I Introduction
The universal character of thermodynamics stems from the type of physical questions it is designed to address. In the thermodynamic limit, experimentally accessible observables probe only coarse-grained averages of the underlying microscopic dynamics, since atomic-scale details lie beyond direct measurement. Such coarse-graining gives rise to the thermodynamic state variables and results in a robust theoretical framework that constrains macroscopic processes largely independently of microscopic details [1]. An essential consequence of this limit is the effective suppression of fluctuations. For small-scale or quantum systems, by contrast, fluctuations play a central role, and under far-from-equilibrium driving the standard thermodynamic description breaks down. In such regimes, one must employ alternative approaches, such as stochastic thermodynamics [2] and information theory [3].
Despite significant advances since Alicki’s pioneering work [4], quantum thermodynamics remains incomplete [5]. Foundational questions persist regarding the definitions of work, heat, and entropy in the quantum regime; see [6, 2] for recent discussions.
@@
Geometrically, gauge theories are expressed in the language of fibre bundles: gauge fields correspond to connections on principal bundles, and their field strengths correspond to the associated curvatures. The base manifold represents spacetime, and the fibres encode the internal symmetry group [7, 8, 9, 10, 11, 8]. Although gauge theories originated in particle physics, they appear in diverse contexts, including low Reynolds number hydrodynamics [12, 13], condensed matter physics [14], quantum information [15], and finance [16], to mention just a few.
@@
Classical thermodynamics has also been explored from this gauge-theoretic perspective. Treating the thermodynamic force as a gauge field, Ref. [17] analysed the Onsager relations within nonequilibrium thermodynamics, while Ref. [18] revealed a connection between $U(1)$ lattice gauge theories and nonequilibrium processes. Moreover, by examining the symmetry under local rescaling of probabilities in Markovian dynamics on finite graphs, Ref. [19] proposed interpreting thermodynamic forces as gauge potentials.
@@
Based on information theory, Refs. [20, 21] cast quantum thermodynamics as a gauge theory of the thermodynamic gauge group. Classical thermodynamics relies on coarse-graining, whereas quantum mechanics assumes full control over microscopic degrees of freedom, yielding an excess of information from a thermodynamic standpoint. The thermodynamic gauge group removes this redundant information, playing a role analogous to the elimination of microscopically irrelevant details in classical thermodynamics. Within this formalism, thermodynamic quantities arise uniquely from gauge invariance, which is one of the most fundamental principles of modern physics.
@@
The gauge structure introduced in [20, 21] is particularly rich, featuring a time-dependent gauge group whose geometric properties remain unexplored. The goal of this work is to clarify the underlying geometry of this theory and to construct the associated fibre-bundle structures in a mathematically rigorous way. This puts thermodynamics on the same mathematical structure as the fundamental physical theories. This may shed new light on the foundations of classical and quantum thermodynamics.
@@
The paper is structured as follows. In Sec. II, we review the gauge theory of the thermodynamics group developed in Refs. [20, 21]. Section III presents how gauge theories can be formulated in terms of the geometry of principal fibre bundles and introduces some common examples from field theory to pave the way for the construction of our theory, which is done in Sec. IV. Finally, in Section V, we discuss how our geometric theory may help in understanding thermodynamics and provide new mathematical insights into the theory by using geometrical and topological techniques, as well as proposing future directions for investigation. Additionally, in Appendix A, we revise the mathematics of fibre bundles, principal bundles, and Lie theory, which are used extensively throughout this paper to make our discussion more self-contained.
@@ II Quantum thermodynamics as a gauge theory
Classical thermodynamics is a theory concerned with the macroscopic bulk properties of large systems, where fluctuations vanish. The theory focusses on the average behaviour of the microscopic degrees of freedom (justified by the central limit theorem), while ignoring the microscopic configurations. This means that, in classical thermodynamics, we have limited access to information, which is referred to as coarse graining [1].
In trying to adapt useful concepts of thermodynamics to the quantum realm, quantum thermodynamics has recently arisen as a powerful theoretical framework [6, 2]. In great contrast to its classical counterpart, however, quantum thermodynamics is usually based on a significant amount of information about the system being studied. For instance, thermodynamic protocols are frequently considered in which knowledge of the density operator is mandatory. This is why quantum thermodynamic quantities are often defined in terms of informational measures [6, 3].
In an attempt to formulate quantum thermodynamics under the same paradigm underlying classical thermodynamics, which encompasses the coarse-graining nature of classical macroscopic measurements, a new framework has been recently proposed [20, 21]. The new theory proposes that when we limit our information about the system to that attainable by measurements of specific observables (thus not performing full state tomography), the full quantum state contains too much information, most of which is inaccessible to the observer. For example, if one is only able to perform energy measurements on the system, it is impossible to distinguish between any two states with the same energy or to obtain information regarding quantum coherences [20, 21].
Figure 1: The coarse graining process defined by restricting the measurements one may perform defines the emergent gauge group of quantum thermodynamics through the removal of redundancies in description due to inaccessible degrees of freedom.
@@
Although we can consider any set of observables, let us consider the case where we can only perform energy measurements on a quantum system described by a time-dependent density matrix $\rho_{t}$. This is the typical situation in many experimental platforms when the dimension of the system increases [22]. Then, for a general time-dependent Hamiltonian $H_{t}$, the mean energy of the system is
@@
with $V_{t}$ representing the symmetry transformation, an element of the thermodynamic group. Since the matrices $V_{t}$ take one density matrix to another, we can assume they are unitary. Joining this fact with Eq. (2), we see that the set of transformations $V_{t}$ allowed must commute with the system’s Hamiltonian. This leads to a gauge group that is isomorphic to the following Lie group [21]:
@@
where $\times$ denotes the Cartesian product and $\mathrm{U}(n_{t}^{i})$ is the unitary group of dimension $n_{t}^{i}$, the (in general, time-dependent) degeneracy of the $i$-th eigenvalue of the Hamiltonian. Thus, $\sum_{i}n_{t}^{i}=d$ is the dimension of the system Hilbert space. The group $\mathrm{G}_{T}$ is called the thermodynamic group. Note that the structure of the group will change if the set of observables is modified; however, the general construction remains the same. Gauge invariance is imposed here by stating that physical quantities, which are expressed as functions of the density matrix, must be invariant under the action of $\mathrm{G}_{T}$.
@@
In further analogy with modern gauge theories, Ref. [20] introduces the notion of a covariant derivative $\mathrm{D}_{t}$ defined by a potential $A_{t}=i\dot{u}_{t}u_{t}^{\dagger}$, where $u_{t}$ is the unitary matrix that diagonalizes the chosen observable, the system’s Hamiltonian in this case. It is defined by the action
@@
In Eq. (8), the $\Pi_{n_{t}^{k}}$ are projectors onto the subspace associated with the degeneracy $n_{t}^{k}$, while $\mathbb{I}_{n_{t}^{k}}$ stands for the identity operator on that subspace. The index $E$ stands for the energy eigenbasis. If we choose a distinct observable, the structure of this operator will remain the same, but the eigenprojectors will be those of the other observable. See Ref. [21] for further details.
Moreover, this procedure for constructing gauge-invariant quantities is not exclusive to entropy. Particularly, since $\rho^{E}$ is the average density matrix with respect to the thermodynamic group (via quantum twirling), it makes sense to define a general gauge-invariant thermodynamic functional as [21]
@@
Thus, this theory defines physical quantities as those invariant under the action of the gauge group $\mathrm{G}_{T}$, in complete analogy to the gauge principle in the fundamental theories of physics. Here, we take a step forward in the formalisation of the theory by providing its constructions in terms of fibre bundles, the natural language in which modern gauge theories are formulated.
@@ III Gauge theories as geometry of principal bundles
In modern mathematics and theoretical physics, fields are understood as sections of fibre bundles. This viewpoint makes explicit the geometric origin of local degrees of freedom, clarifies the role of gauge choices, and highlights how connections generalise ordinary directional derivatives. Fibre bundles, therefore, provide the natural mathematical framework for formulating such theories. For example, in gauge theories describing the fundamental interactions between elementary particles, the interaction is mediated by a gauge potential. Geometrically, a gauge potential is the local expression of a connection that lives on a principal fibre bundle over the four-dimensional spacetime.
@@
Given two topological spaces $\mathds{E}$ (referred to as the total space) and $\mathds{B}$ (referred to as the base space), a Lie group $\mathrm{G}$ (the fibre), and a projection map $\pi\penalty 10000\ :\penalty 10000\ \mathds{E}\penalty 10000\ \rightarrow\penalty 10000\ \mathds{B}$, a principal $\mathrm{G}$-bundle $\xi=(\mathds{E},\pi,\mathrm{G},\mathds{B})$ is a geometric structure such that there exists an open covering $\{O_{\alpha}\}$ and a set of homeomorphisms $\phi_{\alpha}:\pi^{-1}(O_{\alpha})\rightarrow O_{\alpha}\times\mathrm{G}$ (called local trivializations of $\mathds{E}$) satisfying $\pi\phi_{\alpha}^{-1}(x,g)=x$, for $x\in O_{\alpha}$ and $g\in\mathrm{G}$ [10]. Figure 3 presents a schematic representation of these concepts.
@@
Figure 3: A simple example of a bundle. The base space $\mathds{B}$ is the real line $\mathds{R}$, while the total space $\mathds{E}$ is the plane $\mathds{R}\times\mathds{R}$. The projection map $\pi$ associates a fiber $E_{p}$ to each point of $\mathds{B}$. The total space is the union of all the fibres $\mathds{E}=\cup_{p\in\mathds{E}}{E}_{p}$. The projection simply tells you which point of the base space a given fibre element belongs to. In physics, the base space is usually spacetime and the fibre encodes internal degrees of freedom attached to each spacetime point. The section $s$ of the bundle is also shown. A section is a choice of exactly one fibre element over each point of the base space. In physics, a section is the rule that assigns to each point in spacetime the value of a physical quantity living in the fibre.
@@
The above definition reflects the idea that the total space of a principal bundle is locally, but not necessarily globally, a product space of $\mathds{B}$ and $\mathrm{G}$. In the case where $\mathds{E}$ is globally isomorphic to $\mathds{B}\times\mathrm{G}$, we say it is a trivial principal bundle [10]. We can also define a function $s\penalty 10000\ :\penalty 10000\ \mathds{B}\penalty 10000\ \rightarrow\penalty 10000\ \mathds{E}$ called a section of the bundle that locally inverts the projection $\pi$. See Fig. 3 for further details on this definition. A principal bundle is trivial if and only if there exists a global section; i.e., the projection can be globally inverted. The sections of principal fibre bundles in the context of gauge theories are associated with choices of gauge [10, 11].
@@
In the geometric theory of gauge fields, one is generally interested in the following setup [7, 11, 8, 10]: First, in the principal $\mathrm{G}$-bundle $\xi$, the gauge group of the theory is $\mathrm{G}$ (e.g. $\mathrm{U}(1)$ for electromagnetism, $\mathrm{SU}(2)$ for the weak interaction, and $\mathrm{SU}(3)$ for the strong interaction). Most commonly, $\xi$ is the product bundle with $\mathds{B}=\mathcal{M}$, where $\mathcal{M}$ is $(3+1)$-dimensional Minkowski spacetime. This is because $\mathcal{M}$ is effectively $\mathbb{R}^{4}$ with a certain choice of metric; thus, it is contractible [10]. Since $\mathcal{M}$ is contractible, by Theorem 1 in Appendix A, it is always trivial; thus, $\xi$ is isomorphic to the product bundle, which is easier to work with in local coordinates.
@@
Given $\xi$, we also have a vector bundle (a fibre bundle where the fibre has a vector space structure [10, 23]) called the associated bundle. The sections of this bundle are the particles of the theory [11].
The associated bundle of a principal bundle can be constructed in the following manner [23]: given $\xi$ and a linear representation $\rho$ of $\mathrm{G}$ acting on a vector space $\mathds{V}$, the associated bundle $\eta_{\rho}$ is the quotient of $\mathds{E}\times\mathds{V}$ by the following equivalence relation [23]:
@@
where $\star$ is the Hodge star operator. By analysing them component-wise, it can be directly seen that Eqs. (11) leads to all four Maxwell’s equations [7, 10, 8]. This sort of expression is advantageous since it is valid for an arbitrary base manifold and does not depend on any specific set of coordinates [8].
Lastly, it would be of interest to take a look at how charged particle fields arise in this formalism. Following the construction we did earlier for the associated bundle, we must choose a representation of the group $\mathrm{U}(1)$. All of the linear representations of the 1-dimensional unitary group are of the form $\rho(g)=e^{iqt}$, $q\in\mathbb{Z}$ [24]. Thus, choosing a representation is equivalent to choosing an integer. This integer $q$ is called the charge of the particle fields of the associated bundle. Individual matter fields of charge $q$ are then seen to be sections of this associated bundle [11].
The main goal of this work is to formalise the gauge theory of quantum thermodynamics [20, 21] in the language of fibre bundles, thus writing thermodynamics on the same mathematical basis as the fundamental theories of physics.

### v3

# Geometric quantum thermodynamics: A fiber bundle approach

@@ Abstract
Classical thermodynamics is a theory based on coarse-graining, meaning that the thermodynamic variables arise from discarding information related to the microscopic features of the system at hand. In quantum mechanics, however, where one has a high degree of control over microscopic systems, information theory plays an important role in describing the thermal properties of quantum systems. Recently, a new approach has been proposed in the form of quantum thermodynamic gauge theory, where the notion of redundant information arises from a group of physically motivated gauge transformations called the thermodynamic group. In this work, we explore the geometrical structure of quantum thermodynamics. In particular, we do so by explicitly constructing the relevant principal fiber bundle. We then show that there are two distinct (albeit related) geometric structures associated with the gauge theory of quantum thermodynamics. In this way, we express thermodynamics in the same mathematical (geometric) language as the fundamental theories of physics. Finally, we discuss how the geometric and topological properties of these structures may help explain fundamental properties of thermodynamics.
@@ I Introduction
The universal character of thermodynamics stems from the type of physical questions it is designed to address. In the thermodynamic limit, experimentally accessible observables probe only coarse-grained averages of the underlying microscopic dynamics, since atomic-scale details lie beyond direct measurement. This coarse-graining gives rise to thermodynamic state variables and results in a robust theoretical framework that constrains macroscopic processes largely independently of microscopic details [1]. An essential consequence of this limit is the effective suppression of fluctuations. For small-scale or quantum systems, by contrast, fluctuations play a central role, and under far-from-equilibrium driving, the standard thermodynamic description breaks down. In such regimes, alternative approaches must be employed, such as stochastic thermodynamics [2] and information theory [3].
Despite significant advances since Alicki’s pioneering work [4], quantum thermodynamics remains incomplete [5]. Foundational questions persist with respect to the definitions of work, heat, and entropy in the quantum regime; see [6, 2] for recent discussions.
@@
Geometrically, gauge theories are expressed in the language of fiber bundles: gauge fields correspond to connections on principal bundles, and their field strengths correspond to the associated curvatures. The base manifold represents spacetime, and the fibers encode the internal symmetry group [7, 8, 9, 10, 11, 8]. Although gauge theories originated in particle physics, they appear in various contexts, including low Reynolds number hydrodynamics [12, 13], condensed matter physics [14], quantum information [15], and finance [16], to mention just a few.
@@
Classical thermodynamics has also been explored from this gauge-theoretic perspective. Treating the thermodynamic force as a gauge field, Ref. [17] analyzed the Onsager relations within nonequilibrium thermodynamics, while Ref. [18] revealed a connection between $U(1)$ lattice gauge theories and nonequilibrium processes. Moreover, by examining the symmetry under local rescaling of probabilities in Markovian dynamics on finite graphs, Ref. [19] proposed interpreting thermodynamic forces as gauge potentials.
@@
Based on information theory, Refs. [20, 21] cast quantum thermodynamics as a gauge theory of the thermodynamic group. Classical thermodynamics relies on coarse-graining, whereas quantum mechanics assumes full control over microscopic degrees of freedom, yielding an excess of information from a thermodynamic standpoint. The thermodynamic gauge group removes this redundant information, playing a role analogous to the elimination of microscopically irrelevant details in classical thermodynamics. Within this formalism, thermodynamic quantities arise uniquely from gauge invariance, which is one of the most fundamental principles of modern physics.
@@
The gauge structure introduced in [20, 21] is particularly rich, featuring a time-dependent gauge group whose geometric properties remain unexplored. The goal of this work is to clarify the underlying geometry of this theory and to construct the associated fiber-bundle structures in a mathematically rigorous way. This puts thermodynamics on the same mathematical structure as fundamental physical theories. This may shed new light on the foundations of classical and quantum thermodynamics.
@@
The paper is structured as follows. In Sec. II, we review the gauge theory of the thermodynamics group introduced in Refs. [20, 21]. Section III presents how gauge theories can be formulated in terms of the geometry of principal fiber bundles and introduces some common examples from field theory to pave the way for the construction of our theory, which is done in Sec. IV. Finally, in Sec. VI, we discuss how our geometric theory may help us understand thermodynamics and provide new mathematical insights into the theory using geometric and topological techniques, as well as proposing future directions for investigation. Additionally, to make our discussion more self-contained, in Appendix A, we revise the mathematics of fiber bundles, principal bundles, and Lie theory, which are used extensively throughout this paper.
@@ II Quantum thermodynamics as a gauge theory
Classical thermodynamics is a theory concerned with the macroscopic bulk properties of large systems, where fluctuations vanish. The theory focuses on the average behavior of the microscopic degrees of freedom (justified by the central limit theorem), while ignoring the microscopic configurations. This means that, in classical thermodynamics, we have limited access to information, which is called coarse graining [1].
In trying to adapt useful concepts of thermodynamics to the quantum realm, quantum thermodynamics has emerged as a powerful theoretical framework [6, 2]. In great contrast to its classical counterpart, however, quantum thermodynamics is usually based on a significant amount of information about the system being studied. For instance, thermodynamic protocols are frequently considered in which knowledge of the density operator is mandatory. This is why quantum thermodynamic quantities are often defined in terms of informational measures [6, 3].
In an attempt to formulate quantum thermodynamics under the same paradigm that underlying classical thermodynamics, which encompasses the coarse-graining nature of classical macroscopic measurements, a new framework has been proposed [20, 21]. The new theory proposes that when we limit our information about the system to that attainable by measurements of specific observables (thus not performing full state tomography), the full quantum state contains too much information, most of which is inaccessible to the observer. For example, if one can only perform energy measurements on the system, it is impossible to distinguish between any two states with the same energy or to obtain information regarding quantum coherences on this basis [20, 21].
Figure 1: The coarse-graining process defined by restricting the measurements one may perform defines the emergent gauge group of quantum thermodynamics through the removal of redundancies in description due to inaccessible degrees of freedom.
@@
Although we can consider any set of observables, let us consider the case where we can only perform energy measurements on a quantum system described by a time-dependent density matrix $\rho_{t}$. This is the typical situation on many experimental platforms when the dimension of the system increases [22]. Then, for a general time-dependent Hamiltonian $H_{t}$, the mean energy of the system is
@@
with $V_{t}$ representing the symmetry transformation, an element of the thermodynamic group. Since the matrices $V_{t}$ take one density matrix to another, we can assume that they are unitary. Joining this fact with Eq. (2), we see that the set of transformations $V_{t}$ allowed must commute with the Hamiltonian of the system. This leads to a gauge group that is isomorphic to the following Lie group [21]:
@@
where $\times$ denotes the Cartesian product and $\mathrm{U}(n_{t}^{i})$ is the unitary group of dimension $n_{t}^{i}$, the (in general, time-dependent) degeneracy of the $i$-th eigenvalue of the Hamiltonian. Thus, $\sum_{i}n_{t}^{i}=d$ is the dimension of the Hilbert space of the system. The group $\mathrm{G}_{T}$ is called the thermodynamic group. Note that the structure of the group will change if the set of observables is modified; however, the general construction remains the same. Gauge invariance is imposed here by stating that physical quantities, which are expressed as functions of the density matrix, must be invariant under the action of $\mathrm{G}_{T}$.
@@
In general, the degeneracy structure that defines $\mathrm{G}_{T}$ can vary over time, leading to a family of thermodynamic gauge groups $\mathrm{G}_{T}$. The physical interpretation and consequences of this time dependence are discussed in Sec. IV.
In further analogy with modern gauge theories, Ref. [20] introduces the notion of a covariant derivative $\mathrm{D}_{t}$ defined by a potential $A_{t}=i\dot{u}_{t}u_{t}^{\dagger}$, where $u_{t}$ is the unitary matrix that diagonalizes the chosen observable, the Hamiltonian of the system in this case. It is defined by the action
@@
In Eq. (8), $\Pi_{n_{t}^{k}}$ are projectors onto the subspace associated with the degeneracy $n_{t}^{k}$, while $\mathbb{I}_{n_{t}^{k}}$ stands for the identity operator in that subspace. The index $E$ stands for the energy eigenbasis. If we choose a distinct observable, the structure of this operator will remain the same, but the eigenprojectors will be those of the other observable. See Ref. [21] for more details.
Moreover, this procedure for constructing gauge-invariant quantities is not exclusive to entropy. In particular, since $\rho^{E}$ is the average density matrix with respect to the thermodynamic group (via quantum twirling), it makes sense to define a general gauge-invariant thermodynamic functional as [21]
@@
Thus, this theory defines physical quantities as those invariant under the action of the gauge group $\mathrm{G}_{T}$, in complete analogy to the gauge principle in fundamental theories of physics. Here, we take a step forward in the formalization of the theory by providing its constructions in terms of fiber bundles, the natural language in which modern gauge theories are formulated.
@@ III Gauge theories as geometry of principal bundles
In modern mathematics and theoretical physics, fields are understood as sections of fiber bundles. This viewpoint makes explicit the geometric origin of local degrees of freedom, clarifies the role of gauge choices, and highlights how connections generalize ordinary directional derivatives. Fiber bundles, therefore, provide the natural mathematical framework for formulating such theories. For example, in gauge theories describing the fundamental interactions between elementary particles, the interaction is mediated by a gauge potential. Geometrically, a gauge potential is the local expression of a connection that lives on a principal fiber bundle over the four-dimensional spacetime.
@@
Given two topological spaces $\mathds{E}$ (referred to as the total space) and $\mathds{B}$ (referred to as the base space), a Lie group $\mathrm{G}$ (the fiber), and a projection map $\pi\penalty 10000\ :\penalty 10000\ \mathds{E}\penalty 10000\ \rightarrow\penalty 10000\ \mathds{B}$, a principal $\mathrm{G}$-bundle $\xi=(\mathds{E},\pi,\mathrm{G},\mathds{B})$ is a geometric structure such that there exists an open covering $\{O_{\alpha}\}$ and a set of homeomorphisms $\phi_{\alpha}:\pi^{-1}(O_{\alpha})\rightarrow O_{\alpha}\times\mathrm{G}$ (called local trivializations of $\mathds{E}$) satisfying $\pi\phi_{\alpha}^{-1}(x,g)=x$, for $x\in O_{\alpha}$ and $g\in\mathrm{G}$ [10]. Figure 3 presents a schematic representation of these concepts.
@@
Figure 3: A simple example of a bundle. The base space $\mathds{B}$ is the real line $\mathds{R}$, while the total space $\mathds{E}$ is the plane $\mathds{R}\times\mathds{R}$. The projection map $\pi$ associates a fiber $E_{p}$ to each point of $\mathds{B}$. The total space is the union of all the fibers $\mathds{E}=\cup_{p\in\mathds{E}}{E}_{p}$. The projection simply tells you which point of the base space a given fiber element belongs to. In physics, the base space is usually spacetime and the fiber encodes internal degrees of freedom attached to each spacetime point. The section $s$ of the bundle is also shown. A section is a choice of exactly one fiber element over each point of the base space. In physics, a section is the rule that assigns to each point in spacetime the value of a physical quantity living in the fiber.
@@
The above definition reflects the idea that the total space of a principal bundle is locally, but not necessarily globally, a product space of $\mathds{B}$ and $\mathrm{G}$. In the case where $\mathds{E}$ is globally isomorphic to $\mathds{B}\times\mathrm{G}$, we say it is a trivial principal bundle [10]. We can also define a function $s\penalty 10000\ :\penalty 10000\ \mathds{B}\penalty 10000\ \rightarrow\penalty 10000\ \mathds{E}$ called a section of the bundle that locally inverts the projection $\pi$. See Fig. 3 for more details of this definition. A principal bundle is trivial if and only if there exists a global section; i.e., the projection can be globally inverted. The sections of the principal fiber bundles in the context of gauge theories are associated with the choices of gauge [10, 11].
@@
In the geometric theory of gauge fields, one is generally interested in the following setup [7, 11, 8, 10]: First, in the principal $\mathrm{G}$-bundle $\xi$, the gauge group of the theory is $\mathrm{G}$ (e.g. $\mathrm{U}(1)$ for electromagnetism, $\mathrm{SU}(2)$ for the weak interaction, and $\mathrm{SU}(3)$ for the strong interaction). Most often, $\xi$ is the bundle product with $\mathds{B}=\mathcal{M}$, where $\mathcal{M}$ is the Minkowski spacetime of dimensions $(3+1)$. This is because $\mathcal{M}$ is effectively $\mathbb{R}^{4}$ with a certain choice of metric; therefore, it is contractible [10]. Since $\mathcal{M}$ is contractible, according to Theorem 1 in Appendix A, it is always trivial; therefore, $\xi$ is isomorphic to the product bundle, which is easier to work with in local coordinates.
@@
Given $\xi$, we also have a vector bundle (a fiber bundle in which the fiber has a vector space structure [10, 23]) called the associated bundle. The sections of this bundle are the particles of the theory [11].
The associated bundle of a principal bundle can be constructed in the following manner [23]: given $\xi$ and a linear representation $\rho$ of $\mathrm{G}$ acting in a vector space $\mathds{V}$, the associated bundle $\eta_{\rho}$ is the quotient of $\mathds{E}\times\mathds{V}$ by the following equivalence relation [23]:
@@
where $\star$ is the Hodge star operator. By analyzing them component-wise, it can be directly seen that Eqs. (11) lead to all four Maxwell’s equations [7, 10, 8]. This kind of expression is advantageous because it is valid for an arbitrary base manifold and does not depend on any specific set of coordinates [8].
Lastly, it would be of interest to take a look at how charged particle fields arise in this formalism. Following the construction we did earlier for the associated bundle, we must choose a representation of the group $\mathrm{U}(1)$. All linear representations of the unitary 1-dimensional group are of the form $\rho(g)=e^{iqt}$, $q\in\mathbb{Z}$ [24]. Thus, choosing a representation is equivalent to choosing an integer. This integer $q$ is called the charge of the particle fields of the associated bundle. The individual charge fields of matter $q$ are then seen to be sections of this associated bundle [11].
The main goal of this work is to formalize the gauge theory of quantum thermodynamics [20, 21] in the language of fiber bundles, thus writing thermodynamics on the same mathematical basis as the fundamental theories of physics.